decide which is a greater number?
p% of 10 or p% of 20

Answer by stanbon(75887)

decide which is a greater number?
p% of 10 or p% of 20
----
p% = p/100
-----
p% of 10 = (p/100)(10) = (1/10)p
-----
p% of 20 = (p/100)(20) = (1/5)p
-----
When is (1/10)p < (1/5)p ?
Multiply thru by 10 to get:
------
p < 2p
---
The is true as long as p is positive.
----
So p%(20) is greater than p%(10) as long as p is positive.
============================================================
Cheers,
Stan H.